OpenStudy (anonymous):

Find the equation of (3,1)

OpenStudy (goformit100):

@rebeccaskell94

OpenStudy (anonymous):

didnt get that

OpenStudy (anonymous):

I think you are missing something here?

OpenStudy (anonymous):

what I am missing

OpenStudy (anonymous):

would it be y=3x

OpenStudy (anonymous):

@FoolForMath it's so easy...

OpenStudy (anonymous):

help me

OpenStudy (goformit100):

i got it

OpenStudy (anonymous):

what is it

OpenStudy (goformit100):

3y-x=0

OpenStudy (anonymous):

help me solve this

OpenStudy (goformit100):

2y-x=0

OpenStudy (goformit100):

let me explain how

OpenStudy (paxpolaris):

Lines that pass through the origin (0,0), can be written in the general form: \[\huge y= kx\]... where k is the slope, or constant of variation. You can find k by using any point on the line, other than (0,0)\[\Large k=\frac yx\]

OpenStudy (paxpolaris):

using this method 2y-x = 0 would have been written as \[\Large y= \frac 12 x\].... both are right